How many sheets of metal 1/32 inches thick can their be in a pile of 25.5 inches tall?

Answer by checkley71(8403)

25.5/(1/32)=25.5*32=816 SHEETS OF 1/32 IN. METAL IN A PILE 25.5 IN TALL.
PROOF
816(1/32)=816/32=25.5 IN.